Determine the slope of the line that goes through each pair of points.
22. (0,4) and (5,0)

Answer by stanbon(75887)

slope thru (0,4) and (5,0).
--------
slope = (0-4)/(5-0) = -4/5
==================
Cheers,
Stan H.